OpenStudy (2yoututu11):

The number of cars at a dealership dropped from 64 to 48 after a weekend sale. What is the percent of decrease in the number of cars?

OpenStudy (ash2326):

No. of cars dropped from 64 to 48 No. of cars before= 64 No cars after weekend sale= 48 Percentage decrease \[\frac{64-48}{64} \times 100=\frac{16}{64} \times 100\] \[= \frac{\cancel{16} 1}{\cancel{64} 4} \times 100= 25 \%\]

OpenStudy (anonymous):

one percent of 64 = 0.64 one houndred percent of 64 = 64 so 100 times 0.64 = 64 then for 48 cars we want percentage times 0.64 = 48 or percentage = 48 divided by 0.64 then 48 is 75% of 64 and the decrease in percentages is 25